OpenStudy (anonymous):

please help with this question!!!! The cost to mail a package is $5 for the first 2 pounds and 40 cents for each additional ounce. Which of the following functions represents the cost to mail a package if x is the number of ounces over 2 pounds f(x) = 0.4(x + 2) f(x) = 10 + 0.4x f(x) = 5 + 0.4x

OpenStudy (trojanpoem):

f(x) = 10 + 0.4x

OpenStudy (anonymous):

f(x) = 10 + 0.4x

OpenStudy (anonymous):

thank you ! so much so b correct

OpenStudy (anonymous):

dosnt that cross 30,10 on the gragh

OpenStudy (trojanpoem):

The cost = number of items * price We must make all the costs the same dollars or cents I have chosen dollars so 5 dollars , 40 cents / 100 = 0.4 dollar now we don't know the number of additional say they are x so f(X) = number * price ( known) + number * price ( not known) = 5 * 2 + 0.4 * x

OpenStudy (anonymous):

f(x)=10+0.4x

OpenStudy (trojanpoem):

Yeah, but you must understand. How They got it. Or you won't solve any different one.

OpenStudy (anonymous):

i got it ! thanks for the help !

OpenStudy (triciaal):

it should be option C

OpenStudy (triciaal):

$5 for the first 2 pounds